if (1-2-3-4)w=(-1)(-2)(-3)(-4), then w=
(A) -3
(B) -2
(c) -1
(D) 2
(e) 3

Answer by ramkikk66(644)


(1-2-3-4)w = -8w

(-1)(-2)(-3)(-4) = +24 (4 minus signs will give a +ve value for multiplication)

So -8*w = 24

w = 24/(-8) = -3

Ans: A


:)